A student walks and jogs to
college each day. She averages 5km/h walking and 9km/h
jogging. The distance from home to college is
8 km, and she makes the trip in 1 hr. How far does the student log?

Answer by mananth(16949)

A student walks and jogs to
college each day.

and 9 km/h jogging.
let this distance be x
time = x/9
She averages 5km/h walking
this distance will be 8-x
(8-x)/5

she makes the trip in 1 hr.
x/9 + (8-x)/5 =1
multiply by 45
5x+72-9x=45
4x=27
x=27/4
x=6.75 km